# KEK-TH-XXXX

### v1

@@ 1 Introduction
In series of papers [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], Dvali and his collaborators developed the following picture, in which PBHs are effectively stabilized against Hawking evaporation. First, a black hole can be understood as a condensate of a large number of soft gravitons [11, 12, 13]. A crucial assumption is that these soft modes interact with hard modes in such a way that once a critical number of the soft modes is populated, a large number of hard modes gets gapless. This phenomenon is called assisted gaplessness [15, 16, 17, 18]. The soft mode controls the energy level of the hard modes, so the former is also called the master mode. The hard modes can encode an exponential amount of information and are also called memory modes, which are mostly responsible for the entropy of the black hole. A typical microstate of the memory modes contains a large excitation number of gapless modes, whose energy level depends on the excitation number of the master mode. This means that the memory modes backreact on the master mode dynamics since the change of the master mode population would cost huge energy [19, 20]. The slogan is then a black hole, or more generally, any system with large memory modes is effectively stabilized by the burden of its memory. The effect gets important when the black hole loses an $\mathcal{O}(1)$ fraction of its initial mass. Quantitatively, they claimed that this occurs at the latest when the black hole loses half of its initial mass [20]. That is, an older black hole that emitted more than half of its initial mass is not equivalent to a young black hole with the same mass, violating the self-similarity.
@@
After the memory burden effect becomes important, the precise behavior of the black hole is not clear. Dvali et al. proposed either of the following possibilities occur [19, 20]: (i) the Hawking emission rate is significantly suppressed, or (ii) a new classical instability comes in so that the black hole decays into some lumps and gravitational waves. The first possibility opens up a new allowed window for the PBH dark matter mass [19, 20, 22, 23], so we focus on this possibility. [^1] Dark matter can be PBHs in this case but does not have to be so. Dark matter parameter space in the presence of memory-burden PBHs was studied in Refs. [24, 25]. Quantitatively, Dvali et al. argued that the Hawking emission rate is suppressed by some integer power $n_{\text{MB}}$ [defined in Eq. (5) below] of its entropy $S_{\text{BH}}$ [19, 20, 22]. As the black hole entropy is huge, $S_{\text{BH}}\gg 1$, the lifetime of a black hole is significantly extended. However, black holes do evaporate with a suppressed Hawking-emission rate, so we are not considering the absolutely stable remnants of black hole evaporation. [^2] They are in tension with the covariant entropy bound [26]. For an extended discussion on remnants, see Ref. [27]. Induced GWs to probe black hole remnants were studied in Ref. [28].
@@
Cosmological constraints on PBHs beyond the semiclassical regime, i.e., with the memory burden effect were studied in Refs. [22, 23]. In particular, Ref. [23] showed the upper bound on the initial PBH mass of $M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$ from big bang nucleosynthesis (BBN) while the lower bound from galactic and extragalactic gamma rays [29, 4] depends on the suppression power $n_{\text{MB}}$ of the memory burden effect. For $n_{\text{MB}}=1$, the allowed region is almost closed. For $n_{\text{MB}}=2$, the allowed mass is $10^{5}\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$. While $n_{\text{MB}}>2$ allows lighter PBHs and this extension is straightforward, the associated induced GWs have too high frequencies to be detected in the near future. Assuming the post-inflationary PBH formation mechanism in the radiation-dominated era [30, 31], the lower bound on the PBH mass is $M_{\text{PBH,ini}}\gtrsim 0.5\,\text{g}$, which comes from the upper bound on the tensor-to-scalar ratio $r\lesssim 0.04$ [32, 33, 34] with the measured amplitude of the primordial curvature perturbations $A_{\text{s}}=2.1\times 10^{-9}$ [35, 36]. Therefore, we consider the mass range $0.5\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$ but focus on $10^{5}\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$.
@@
In this paper, we study the induced GWs associated with the PBH dark matter in the presence of the memory burden effect. Our focus is the simple scenario in which the induced GWs are produced around the time of PBH formation in the radiation-dominated era. [^3] Recently, another induced-GW signal related to the non-dark-matter PBHs was studied in Ref. [47, 25, 48]. In their scenario, PBHs dominate the Universe and then evaporate to reheat the Universe. If the distributions of the mass and spin of PBHs are negligible, the PBHs evaporate almost simultaneously leading to a rapid change of the equation of state of the Universe. In such an occasion, the induced GWs are significantly enhanced by the poltergeist mechanism [49] (named in Ref. [50]), where the perturbations (of the PBH number density [50]) were assumed to be adiabatic. It was generalized to the Poissonian case [51] in Ref. [52]. The Poissonian contribution was considered in Refs. [47, 25], while both contributions were considered in Ref. [48]. See also Refs. [53, 50, 54, 55, 56] for the importance of the finite transition time scale, which makes the enhancement of the induced GWs milder. Note also that the mass distribution of the memory-burden PBHs was discussed in Ref. [21], which suppresses the induced GWs from the poltergeist mechanism [50, 54]. In section 2, we review PBHs discussing their formation, evaporation, and abundance in the presence of the memory burden effect. We introduce example power spectra of the primordial curvature perturbations. Section 3 is the main section of this paper, where we discuss the induced-GW signals of the memory-burdened PBH dark matter scenario. We discuss the observational prospects of the induced GWs. In Section 4, we discuss some possibilities to cross-check or exclude the memory-burdened PBH dark matter scenario. For example, we argue that the memory burden effect can be excluded if the induced GWs associated with the light PBHs are too strong. We also discuss another GW signal, which is from mergers of PBHs, to confirm the memory burden effect although the frequency of the GWs is extremely high. We summarize the paper and conclude in section 5. We use the natural unit $c=\hbar=k_{\text{B}}=8\pi G=1$ unless explicitly denoted.
@@ 2 Primordial black hole abundance with memory burden
Although the PBH abundance is sensitive to the details of the calculation method, [^4] See Refs. [57, 58, 59, 60, 61, 62] for recent discussions on (uncertainties of) the PBH abundance. For more general reviews of PBHs, see Refs. [63, 4, 5, 7]. the difference can be absorbed by the amplitude of the curvature perturbations. Since the PBH abundance is extremely sensitive to it while the intensity of the induced GWs, which is the main topic of this paper, is only quadratically sensitive, our results do not depend on the details of the calculation formalism of the PBH abundance. Therefore, we take a simple method; we basically follow the method in Refs. [64, 57, 65, 66]. [^5] The differences are as follows. We use $\delta_{\text{c}}=0.42$ rather than $0.45$. This means that PBHs are more easily produced. To adjust the PBH abundance, a smaller amplitude of the curvature perturbations is required. This leads to a conservative estimate for the induced GWs. We also use $\gamma=(1/\sqrt{3})^{3}$ rather than $0.2$ following Ref. [67].
@@
$\dot{M}_{\text{PBH}}=$ $\begin{cases}-\frac{\pi\mathcal{G}g_{*\text{H}}}{480M_{\text{PBH}}^{2}}&(M_{\text{PBH}}\geq qM_{\text{PBH,ini}})\\ -\frac{\pi\mathcal{G}g_{*\text{H}}}{480S_{\text{BH}}^{n_{\text{MB}}}M_{\text{PBH}}^{2}}&(M_{\text{PBH}}<qM_{\text{PBH,ini}})\end{cases}$ (5)
where $\mathcal{G}\simeq 3.8$ is the gray-body factor, $g_{*\text{H}}(\simeq 108$ for $M_{\text{PBH}}\ll 10^{11}\text{g}$ in the Standard Model [69] at least in the conventional case without memory burden) is the effective number of degrees of freedom emitted by Hawking radiation, and $1/2\leq q\leq 1$ parameterizes the critical mass at which the memory burden effect becomes relevant. The Bekenstein-Hawking entropy of the black hole is given by $S_{\text{BH}}(M_{\text{PBH}})=2\pi A=M_{\text{PBH}}^{2}/2$ [72, 2] in the reduced Planck unit with $A=M_{\text{PBH}}^{2}/(4\pi)$ being the surface area. There are some caveats in the above expression. First, the detailed form of the emission rate in the memory-burden regime is unknown. Beyond the semiclassical regime, it is not clear if we can rely on notions of the black-hole thermodynamics [73] such as Hawking temperature and entropy. Second, the above evaporation rate is discontinuous at $M_{\text{PBH}}=qM_{\text{PBH,ini}}$. Realistically, the evaporation rate will smoothly transform from the semiclassical regime to the memory burden regime because the magnitude of the backreaction from the memory modes to the master mode is continuous. Third, there are several proposals [22, 23, 47] for the detailed form of the evaporation rate in the memory burden regime. For example, some authors replace $M_{\text{PBH}}$ and/or $S_{\text{BH}}(M_{\text{PBH}})$ with $qM_{\text{PBH,ini}}$ and/or $S_{\text{BH}}(qM_{\text{PBH,ini}})$. These differences lead to only negligible effects on observables at present at least in the parameter range of our main interest $10^{5}\leq M_{\text{PBH,ini}}/\mathrm{g}\leq 10^{10}$. For $n_{\text{MB}}\geq 2$, it is a good approximation to set $M_{\text{PBH}}|_{\text{present}}=qM_{\text{PBH,ini}}$ since the Hawking emission rate in the semiclassical regime for these masses is so effective while the evaporation rate is significantly suppressed in the memory burden regime.
@@ 3 Induced gravitational waves
where $\eta$ is the conformal time, $h=H_{0}/(100\,\mathrm{km}/\mathrm{s}/\mathrm{Mpc})$ is the reduced Hubble parameter, $k=2\pi f$ is the wavenumber of the GWs, $\Omega_{\text{GW}}(\eta_{\text{c}})$ is the value of $\Omega_{\text{GW}}(\eta)$ that reached a constant value during the radiation-dominated epoch, $u=(t+s+1)/2$, v=(ts+1)/2v=(t-s+1)/2, and the oscillation average (denoted by the overline) of the kernel function $\lim_{\eta\to\infty}\overline{(k\eta)^{2}I(t,s,k\eta)^{2}}$ is [76, 75]
@@
[…]
Figure 1: The spectra of the induced GWs are shown in monochrome colors. From left (black) to right (light gray), $M_{\text{PBH,ini}}/\mathrm{g}=10^{10}$, $10^{9}$, $10^{8}$, $10^{7}$, $10^{6}$, and $10^{5}$. The left and right panels show the cases with $\mathcal{P}_{\zeta}$ being the delta function and the lognormal function with $\Delta=1$, respectively. The colored lines denote the power-law integrated sensitivity curves for 1-year observations adopted from Ref. [77]. The observations include LISA (sky blue), DECIGO (bluish green), BBO (vermilion), ET (blue), CE (orange), HLVK (design sensitivity) (reddish purple), and HLV O3 (yellow shaded).
The spectra of the induced GWs are plotted in Fig. 1 with the power-law integrated sensitivity curves of various GW observations assuming $1$ year observations. The left panel shows the case of the delta-function $\mathcal{P}_{\zeta}(k)$, and the right panel shows the case of the lognormal $\mathcal{P}_{\zeta}(k)$. Since a small-scale mode re-enters the Hubble horizon earlier, the induced GWs associated with a smaller PBH have higher frequencies.
The power-law integrated sensitivity curves in Fig. 1 are those of the following GW observations. The sky blue, bluish green, and vermilion lines correspond to LISA [45, 46], DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) [81, 82, 83, 84], and Big Bang Observer (BBO) [85, 86, 87]. These are space-based interferometers. The shaded yellow region is the Laser Interferometer Gravitational-Wave Observatory (LIGO)/Virgo (HLV) O3 constraint on SGWB [88]. This is an existing constraint whereas the other lines are sensitivity curves for future observations taken from Ref. [77]. The pink line corresponds to a combination of advanced LIGO [89, 90], advanced Virgo [91], and Kamioka Gravitational wave detector (KAGRA) [92, 93, 94, 95, 96] (HLVK). These are ground-based interferometers. The next generation ground-based interferometers include Einstein Telescope (ET) [97, 98, 99, 100] (the blue line) and Cosmic Explorer (CE) [101, 102] (the orange line).
[…]
Figure 2: The signal-to-noise ratio (SNR) for 1-year observations. Left: the SNR as a function of the initial PBH mass. The solid and dashed lines correspond to the cases with $\mathcal{P}_{\zeta}(\ln k)$ being the Gaussian with $\Delta=1$ or the delta function, respectively. The orange, blue, vermilion, bluish-green, and reddish-purple lines correspond to CE, ET, BBO, DECIGO, and HLVK, respectively. Right: the contour plot of $\log_{10}\text{SNR}$ for CE (1 year) on the plane of $\log_{10}(M_{\text{PBH,ini}}/\text{g})$ and $\log_{10}\Delta$. The contour lines denote 4-4, 3-3, …, 33 from left to right; the border between the pinkish and greenish regions corresponds to $\text{SNR}=1$. See Appendix B for similar plots for ET, BBO, DECIGO, and HLVK.
$\text{SNR}=$ $\sqrt{n_{\text{det}}t_{\text{obs}}\int_{f_{\text{min}}}^{f_{\text{max}}}\mathrm{d}f\left(\frac{\Omega_{\text{signal}}(f)}{\Omega_{\text{noise}}(f)}\right)^{2}},$ (10)

### v2

# KEK-TH-2654 KEK-Cosmo-0358 KEK-QUP-2024-0021

@@ 1 Introduction
In series of papers [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], Dvali and his collaborators developed the following picture, in which PBHs are effectively stabilized against Hawking evaporation. First, a black hole can be understood as a condensate of a large number of soft gravitons [11, 12, 13]. A crucial assumption is that these soft modes interact with hard modes in such a way that once a critical number of the soft modes is populated, a large number of hard modes gets gapless. This phenomenon is called assisted gaplessness [15, 16, 17, 18]. The soft mode controls the energy level of the hard modes, so the former is also called the master mode. The hard modes can encode an exponential amount of information and are also called memory modes, which are mostly responsible for the entropy of the black hole. A typical microstate of the memory modes contains a large excitation number of gapless modes, whose energy levels depend on the occupation number of the master mode. This means that the memory modes backreact on the master mode dynamics since the change of the master mode population would cost huge energy [19, 20]. Then, the slogan is that a black hole, or more generally, any system with large memory modes is effectively stabilized by the burden of its memory. The effect gets important at the latest when the black hole loses an $\mathcal{O}(1)$ fraction of its initial mass. [^1] It may become important as soon as an $\mathcal{O}(1/\sqrt{S_{\text{BH}}})$ fraction of its initial mass is emitted [22, 23] where $S_{\text{BH}}$ is the initial entropy of the black hole. Quantitatively, they claimed that this occurs at the latest when the black hole loses half of its initial mass [20]. That is, an older black hole that emitted more than half of its initial mass is not equivalent to a young black hole with the same mass, violating the self-similarity.
@@
After the memory burden effect becomes important, the precise behavior of the black hole is not clear. Dvali et al. proposed either of the following possibilities occur [19, 20]: (i) the Hawking emission rate is significantly suppressed, or (ii) a new classical instability comes in so that the black hole decays into some lumps and gravitational waves. The first possibility opens up a new allowed window for the PBH dark matter mass [19, 20, 24, 25], so we focus on this possibility.222 Dark matter can be PBHs in this case but does not have to be so. Dark matter parameter space in the presence of memory-burden PBHs was studied in Refs. [26, 27]. Quantitatively, Dvali et al. argued that the Hawking emission rate is suppressed by some integer power $n_{\text{MB}}$ [defined in Eq. (5) below] of its entropy $S_{\text{BH}}$ [19, 20, 24]. As the black hole entropy is huge, $S_{\text{BH}}\gg 1$, the lifetime of a black hole is significantly extended. However, black holes do evaporate with a suppressed Hawking-emission rate, so we are not considering the absolutely stable remnants of black hole evaporation.333 They are in tension with the covariant entropy bound [28]. For an extended discussion on remnants, see Ref. [29]. Induced gravitational waves to probe black hole remnants were studied in Refs. [30, 31].
Figure 1: Constraints on the PBH parameter space with the memory burden effect adopted from Ref. [25]. Left: the blue solid, orange dashed, and green dotted lines are the lower bounds on the memory-burden parameter $n_{\text{MB}}$ from the galactic gamma-ray background, extragalactic gamma-ray background, and cosmic microwave background (CMB), respectively. In the black shaded region, PBHs have evaporated by the present time even with the memory burden effect. The range $10^{10}\,\mathrm{g}\leq M_{\text{PBH,ini}}\leq 10^{17}\,\mathrm{g}$ is shaded in vermilion to approximately show the constraints from the BBN. Right: the upper bound on the PBH abundance $f_{\text{PBH}}$ in the case of $n_{\text{MB}}=2$. The blue solid, orange dashed, green dotted, and vermilion dot-dashed lines are the constraints from the galactic gamma-ray background, extragalactic gamma-ray background, CMB, and BBN, respectively. To the left of the black vertical line, the PBHs have evaporated by the present time. See Ref. [25] for more details.
@@
Cosmological constraints on PBHs beyond the semiclassical regime, i.e., with the memory burden effect were studied in Refs. [24, 25]. In particular, Ref. [25] showed the upper bound on the initial PBH mass of $M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$ from big bang nucleosynthesis (BBN) while the lower bound from galactic and extragalactic gamma rays [32, 4] depends on the suppression power $n_{\text{MB}}$ of the memory burden effect (see the left panel of Fig. 1). For $n_{\text{MB}}=1$, the allowed region is almost closed. For $n_{\text{MB}}=2$, the allowed mass is $10^{5}\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$ (see the right as well as left panel of Fig. 1). While $n_{\text{MB}}>2$ allows lighter PBHs and this extension is straightforward, the associated induced gravitational waves (GWs) (introduced in the next paragraph) have too high frequencies to be detected in the near future. Assuming the post-inflationary PBH formation mechanism in the radiation-dominated era [33, 34], the lower bound on the PBH mass is $M_{\text{PBH,ini}}\gtrsim 0.5\,\text{g}$, which comes from the upper bound on the tensor-to-scalar ratio $r\lesssim 0.04$ [35, 36, 37] with the measured amplitude of the primordial curvature perturbations $A_{\text{s}}=2.1\times 10^{-9}$ [38, 39]. Therefore, we consider the mass range $0.5\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$ but focus on $10^{5}\,\mathrm{g}\lesssim M_{\text{PBH,ini}}\lesssim 10^{10}\,\mathrm{g}$.
@@
In this paper, we study the induced GWs associated with the PBH dark matter in the presence of the memory burden effect. [^4] GWs from Hawking radiation of PBHs in the presence of the memory burden effect were studied in Ref. [50], which appeared on arXiv simultaneously with this paper. Our focus is the simple scenario in which the induced GWs are produced around the time of PBH formation in the radiation-dominated era.555 Recently, another induced-GW signal related to the non-dark-matter PBHs was studied in Refs. [51, 27, 52]. In their scenario, PBHs dominate the Universe (see also Ref. [53] for the GWs induced by PBH isocurvature perturbations) and then evaporate to reheat the Universe. If the distributions of the mass and spin of PBHs are negligible, the PBHs evaporate almost simultaneously leading to a rapid change of the equation of state of the Universe. In such an occasion, the induced GWs are significantly enhanced by the poltergeist mechanism [54] (named in Ref. [55]), where the perturbations (of the PBH number density [55]) were assumed to be adiabatic. It was generalized to the Poissonian case [56] in Ref. [57]. The Poissonian contribution was considered in Refs. [51, 27], while both contributions were considered in Ref. [52]. See also Refs. [58, 55, 59, 60, 61] for the importance of the finite transition timescale, which makes the enhancement of the induced GWs milder. Note also that the mass distribution of the memory-burden PBHs was discussed in Ref. [21], which suppresses the induced GWs from the poltergeist mechanism [55, 59]. In Sec. 2, we review PBHs discussing their formation, evaporation, and abundance in the presence of the memory burden effect. We introduce example power spectra of the primordial curvature perturbations. Sec. 3 is the main section of this paper, where we discuss the induced-GW signals of the memory-burdened PBH dark matter scenario. We discuss the observational prospects of the induced GWs. In Sec. 4, we discuss some possibilities to cross-check or exclude the memory-burdened PBH dark matter scenario. For example, we argue that the memory burden effect can be excluded if the induced GWs associated with the light PBHs are too strong. We also discuss another GW signal, which is from mergers of PBHs, to confirm the memory burden effect although the frequency of the GWs is extremely high. We summarize the paper and conclude in Sec. 5. We use the natural unit $c=\hbar=k_{\text{B}}=8\pi G=1$ unless explicitly denoted.
@@ 2 Primordial black hole abundance with memory burden
Although the PBH abundance is sensitive to the details of the calculation method,666 See Refs. [62, 63, 64, 65, 66, 67] for recent discussions on (uncertainties of) the PBH abundance. For more general reviews of PBHs, see Refs. [68, 4, 5, 7]. the overall difference can be absorbed by the amplitude of the curvature perturbations. Since the PBH abundance is extremely sensitive to it while the intensity of the induced GWs, which is the main topic of this paper, is only quadratically sensitive, our results do not depend on the details of the calculation formalism of the PBH abundance. Therefore, we take a simple method; we basically follow the method in Refs. [69, 62, 70, 71].777 The differences are as follows. We use $\delta_{\text{c}}=0.42$ rather than $0.45$. This means that PBHs are more easily produced. To adjust the PBH abundance, a smaller amplitude of the curvature perturbations is required. This leads to a conservative estimate for the induced GWs. We also use $\gamma=(1/\sqrt{3})^{3}$ rather than $0.2$ following Ref. [72].
@@
$\dot{M}_{\text{PBH}}=$ $-\frac{\pi\mathcal{G}g_{*\text{H}}}{480M_{\text{PBH}}^{2}}\times\begin{cases}1&(M_{\text{PBH}}\geq qM_{\text{PBH,ini}})\\ S_{\text{BH}}^{-n_{\text{MB}}}&(M_{\text{PBH}}<qM_{\text{PBH,ini}})\end{cases}$ (5)
where $\mathcal{G}\simeq 3.8$ is the gray-body factor, $g_{*\text{H}}(\simeq 108$ for $M_{\text{PBH}}\ll 10^{11}\text{g}$ in the Standard Model [74] at least in the conventional case without memory burden) is the effective number of degrees of freedom emitted by Hawking radiation, and $1/2\leq q\leq 1$ parametrizes the critical mass at which the memory burden effect becomes relevant. The Bekenstein-Hawking entropy of the black hole is given by $S_{\text{BH}}(M_{\text{PBH}})=2\pi A=M_{\text{PBH}}^{2}/2$ [77, 2] in the reduced Planck unit with $A=M_{\text{PBH}}^{2}/(4\pi)$ being the surface area. There are some caveats in the above expression. First, the detailed form of the emission rate in the memory-burden regime is unknown. Beyond the semiclassical regime, it is not clear if we can rely on notions of the black-hole thermodynamics [78] such as Hawking temperature and entropy. Second, the above evaporation rate is discontinuous at $M_{\text{PBH}}=qM_{\text{PBH,ini}}$. Realistically, the evaporation rate will smoothly transform from the semiclassical regime to the memory burden regime because the magnitude of the backreaction from the memory modes to the master mode is continuous. Third, there are several proposals [24, 25, 51] for the detailed form of the evaporation rate in the memory burden regime. For example, some authors replace $M_{\text{PBH}}$ and/or $S_{\text{BH}}(M_{\text{PBH}})$ with $qM_{\text{PBH,ini}}$ and/or $S_{\text{BH}}(qM_{\text{PBH,ini}})$. These differences lead to only negligible effects on observables at present at least in the parameter range of our main interest $10^{5}\leq M_{\text{PBH,ini}}/\mathrm{g}\leq 10^{10}$. For $n_{\text{MB}}\geq 2$, it is a good approximation to set $M_{\text{PBH}}=qM_{\text{PBH,ini}}$ at present since the Hawking emission rate in the semiclassical regime for these masses is so effective while the evaporation rate is significantly suppressed in the memory burden regime.
@@ 3 Induced gravitational waves
where $\eta$ is the conformal time, $h=H_{0}/(100\,\mathrm{km}/\mathrm{s}/\mathrm{Mpc})$ is the reduced Hubble parameter, $k=2\pi f$ is the wavenumber of the GWs, $\Omega_{\text{GW}}(\eta_{\text{c}})$ is the value of $\Omega_{\text{GW}}(\eta)$ that reached a constant value during the radiation-dominated epoch, $u=(t+s+1)/2$, $v=(t-s+1)/2$, and the oscillation average (denoted by the overline) of the kernel function $\lim_{\eta\to\infty}\overline{(k\eta)^{2}I(t,s,k\eta)^{2}}$ is [81, 80]
@@
Figure 2: The spectra of the induced GWs associated with the PBH dark matter are shown in monochrome colors. From left (black) to right (light gray), PBHs with $2M_{\text{PBH}}/\mathrm{g}=M_{\text{PBH,ini}}/\mathrm{g}=10^{10}$, $10^{8}$, $10^{6}$, $10^{4}$, $10^{2}$, and $10^{0}$ are assumed to comprise the totality of dark matter with the help of the memory burden effect. The left and right panels show the cases with $\mathcal{P}_{\zeta}$ being the delta function and the lognormal function with $\Delta=1$, respectively. The colored lines in the region $f<10^{4}\,\mathrm{Hz}$ denote the power-law integrated sensitivity curves for 1-year observations adopted from Ref. [82]. The observations include LISA (sky blue), DECIGO (bluish green), BBO (vermilion), ET (blue), CE (orange), HLVK (design sensitivity) (reddish purple), and HLV O3 (yellow shaded). In addition, the sensitivity of future resonant cavities [83, 84], assuming the detectability of the strain of $10^{-14}\,\mathrm{W}$, is adopted from Ref. [85] and shown by the thin purple line.
The spectra of the induced GWs are plotted in Fig. 2 with various observational sensitivity curves. The left panel shows the case of the delta-function $\mathcal{P}_{\zeta}(k)$, and the right panel shows the case of the lognormal $\mathcal{P}_{\zeta}(k)$. Since a small-scale mode re-enters the Hubble horizon earlier, the induced GWs associated with a smaller PBH have higher frequencies. For simplicity, we set the present PBH mass to be the half of its initial value $M_{\text{PBH}}=M_{\text{PBH,ini}}/2$ for all choices of $M_{\text{PBH,ini}}$ in the figure.
The colored lines in the region $f<10^{4}\,\mathrm{Hz}$ in Fig. 2 are the power-law integrated sensitivity curves of the following GW observations assuming 1-year observations. The sky blue, bluish green, and vermilion lines correspond to LISA [48, 49], DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) [89, 90, 91, 92], and Big Bang Observer (BBO) [93, 94, 95]. These are space-based interferometers. The shaded yellow region is the Laser Interferometer Gravitational-Wave Observatory (LIGO)/Virgo (HLV) O3 constraint on SGWB [96]. This is an existing constraint whereas the other lines are sensitivity curves for future observations taken from Ref. [82]. The pink line corresponds to a combination of advanced LIGO [97, 98], advanced Virgo [99], and Kamioka Gravitational wave detector (KAGRA) [100, 101, 102, 103, 104] (HLVK). These are ground-based interferometers. The next generation ground-based interferometers include Einstein Telescope (ET) [105, 106, 107, 108] (the blue line) and Cosmic Explorer (CE) [109, 110] (the orange line). For the reference of high-frequency GW observations, we also plot the sensitivity [84, 85] of futuristic resonant cavities [83, 84] by the thin purple line in the region $f>10^{5}\,\mathrm{Hz}$, assuming the detectability of the strain of $10^{-14}\,\mathrm{W}$.
Figure 3: The signal-to-noise ratio (SNR) for 1-year observations. Left: the SNR as a function of the initial PBH mass. The solid and dashed lines correspond to the cases with $\mathcal{P}_{\zeta}(\ln k)$ being the Gaussian with $\Delta=1$ or the delta function, respectively. The orange, blue, vermilion, bluish-green, and reddish-purple lines correspond to CE, ET, BBO, DECIGO, and HLVK, respectively. Right: the contour plot of $\log_{10}\text{SNR}$ for CE (1 year) on the plane of $\log_{10}(M_{\text{PBH,ini}}/\text{g})$ and $\log_{10}\Delta$. The contour lines denote $-4$, $-3$, …, $3$ from left to right; the border between the pinkish and greenish regions corresponds to $\text{SNR}=1$. See Appendix B for similar plots for ET, BBO, DECIGO, and HLVK.